# Rationale-based Ensemble of Multiple QA Strategies for Zero-shot Knowledge-based VQA

### v2

@@ 4.3 Ablation Studies
Analysis of RSF.
We incrementally integrate automatic and mechanistic rationales into REACT. The results are shown in Tab. 3. The additional baseline (#1) selects answer candidate randomly without LLM-based answer fusion. The ablation #2 selects answers using an LLM without rationales. #5 replaces the mechanistic rationale with a random sentence from the decision context.
Comparing #1 with #2, we observe minor improvements (0.6% and 1.1%) from the simplistic answer ensemble without rationales. #3 and #4 improve over #2 by 0.5% and 0.7% on OKVQA and by 0.3% and 1.1% on A-OKVQA, demonstrating both types of rationales are effective by themselves. Adding a random mechanistic rationale to #4, we obtain ablation #5, which retains the same performance on OKVQA but decreases by 0.6% on A-OKVQA. This highlights the importance of selecting proper mechanistic rationales, as wrong rationales are harmful. Most interestingly, leveraging both rationales, #6 attains the best performance, showing the two rationales are complementary.
Finally, with Fig. 5, we analyze the relation between the rationale surface form containing exactly the answer candidate, which we refer to as a “hit”, and the candidate being selected by the RSF LLM. Similar to the finding of Tiong et al. (2022), we observe a positive correlation between the two events. This suggests that the surface form of the rationales plays a role in the decisions of the answer-selecting LLM. The largest differential is with caption-only, where having the answer in either rationale increases the selection probability by 22.5% [^2] P(selected $\mid$ either hits)--P(selected $\mid$ neither hits)=22.5%. In contrast, having the candidate in the rationales under SK or LK is not as important (probability differences of 17.5-10.3%). A possible reason is that the LLM conducts more complex operations on questions requiring external knowledge than simple surface-form matching.
@@ 4.4 Case Study
We show some example K-VQA questions in Fig. 6. To conserve space, please consult Fig. 1 for the corresponding captions and knowledge details. In strategy ensemble, we provide answer candidates with corresponding rationales and caption-generated QA pairs in the prompt. More examples can be found in Tab. 7 of the Appendix.
In example (a), since “Brick walkway” appears in the captions, the answer from caption-only strategy is “Brick”. However, as concrete is a more common street material, the answers from short-form and long-form knowledge are both “Concrete”. Still, the model selects "Brick" as the final answer, possibly due to the strength of the rationale. In example (b), the answer from the caption-only strategy is "Muffin", but the its rationales are not proper explanations. The answer from the short-form knowledge strategy is corect and has sensible rationales, and it is selected as the final answer. In example (c), desert and savanna are both possible answers, which are hard to distinguish without visual information. The model may have been informed by the rationales from the second strategy mentioning rivers, which are more common in a savanna than in a desert.
@@ 5 Conclusion
This paper proposes Rationale-based Ensemble of Answer Context Tactics (REACT) to achieve dynamic ensemble of multiple question-answering tactics. REACT first generates three types of decision contexts: image captions, short-form knowledge and long-form knowledge, as well as answers generated from each decision context separately. After that, we generate two types of rationales for each answer. An LLM selects the best answer from the rationales. REACT achieves state-of-the-art results on OK-VQA among comparable methods and showcases LLM-informed ensemble as an effective VQA strategy.
@@
In this paper, we generate three different types of decision contexts for each question, including captions, short-form knowledge, and long-form knowledge, where both types of knowledge are generated based on captions. However, in the process of converting images to captions, some visual details are inevitably lost, resulting in inaccurate captions. Therefore, the knowledge generated based on these captions may also be inaccurate. These inaccurate decision contexts may lead to incorrect answers. In extreme cases, if all answer candidates are wrong, our method will fail. Hence, improving the quality of decision contexts is an important direction for future research. Moreover, our method requires multiple calls to the LLM for inference, which is time-consuming. Reducing the running time would be a direction for future research.
Although our work has achieved good results, it also inherits the social-economic biases of LLMs and datasets about gender, age, race, and ethnicity. Future work can focus on addressing these issues.
Appendix A Prompt Templates
@@
Table 7: Prompt template for strategy fusion. We first provide rationales for each answer, then ask the model to select a final answer from the answer candidates based on these rationales.
Appendix B Examples
In this section, we display more examples from OK-VQA where REACT can help model select the correct answer, as shown in Fig 7.
Figure 7: Examples from OKVQA where our method can help model select the correct answer. From top to bottom are the caption-only strategy, the short-form knowledge strategy, and the long-form knowledge strategy.

### v4

@@ 4.3 Ablation Studies
The Rationalization Strategies.
We incrementally integrate automatic and mechanistic rationales into DietCoke. The results are shown in Tab. 3. The additional baseline (#1) selects answer candidate randomly without LLM-based answer fusion. The ablation #2 selects answers using an LLM without rationales. #5 replaces the mechanistic rationale with a random sentence from the decision context.
Comparing #1 with #2, we observe minor improvements (0.6% and 1.1%) from the simplistic answer ensemble without rationales. #3 and #4 improve over #2 by 0.5% and 0.7% on OKVQA and by 0.3% and 1.1% on A-OKVQA, demonstrating both types of rationales are effective by themselves. Adding a random mechanistic rationale to #4, we obtain ablation #5, which retains the same performance on OKVQA but decreases by 0.6% on A-OKVQA. This highlights the importance of selecting proper mechanistic rationales, as wrong rationales are harmful. Most interestingly, leveraging both rationales, #6 attains the best performance, +1.2% / +1.4% over #2. Notably, the improvements are exactly the sum of the improvements attained by the individual rationales (#3 and #4), suggesting the two rationales are perfectly complementary and the ensemble over rationales is effective.
Finally, with Fig. 5, we analyze the relation between the rationale surface form containing exactly the answer candidate, which we refer to as a “hit”, and the candidate being selected by the ensemble phase LLM. Similar to the finding of Tiong et al. (2022), we observe a positive correlation between the two events. This suggests that the surface form of the rationales plays a role in the decisions of the answer-selecting LLM. The largest differential is with caption-only, where having the answer in either rationale increases the selection probability by 22.5% [^2] P(selected $\mid$ either hits)--P(selected $\mid$ neither hits)=22.5%. In contrast, having the candidate in the rationales under SK or LK is not as important (probability differences of 17.5-10.3%). A possible reason is that the LLM conducts more complex operations on questions requiring external knowledge than simple surface-form matching.
@@ 4.4 Case Study
We show some example K-VQA questions in Fig. 6. To conserve space, we do not repeat the corresponding captions and knowledge details, available in Fig. 1. In strategy ensemble, we provide answer candidates with corresponding rationales and caption-generated QA pairs in the prompt. More examples can be found in Fig. 7 of the Appendix B.
In example (a), since “Brick walkway” appears in the captions, the answer from caption-only strategy is “Brick”. However, as concrete is a more common street material, the answers from short-form and long-form knowledge are both “Concrete”. Still, the model selects "Brick" as the final answer, possibly due to the strength of the rationale. In example (b), the answer from the caption-only strategy is "Muffin", but the its rationales are not proper explanations. The answer from the short-form knowledge strategy is correct and has sensible rationales, and it is selected as the final answer. In example (c), desert and savanna are both possible answers, which are hard to distinguish without visual information. The model may have been informed by the rationales from the second strategy mentioning rivers, which are more common in a savanna than in a desert.
@@ 5 Conclusion
We propose DietCoke, which ensembles several QA strategies for knowledge-based VQA. DietCoke first generates three types of decision contexts: image captions, short-form knowledge and long-form knowledge, and answers the question from each decision context separately. After that, DietCoke generates two types of rationales for each answer. An ensemble LLM selects the best answer from the rationales. DietCoke achieves state-of-the-art results on OK-VQA among comparable methods and showcases LLM-informed, rationale-based ensemble as an effective VQA method. DietCoke allows potential trade-off between compute and answer quality and may contribute to research on scaling LLMs with inference-time compute, a research direction pioneered by OpenAI-o1 OpenAI (2024).
Acknowledgments
We gratefully acknowledge the support by the Nanyang Associate Professorship and the National Research Foundation Fellowship (NRF-NRFF13-2021-0006), Singapore. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not reflect the views of the funding agencies.
@@ Acknowledgments
In this paper, we generate three different types of decision contexts for each question, including captions, short-form knowledge, and long-form knowledge, where both types of knowledge are generated based on captions. However, in the process of converting images to captions, some visual details are inevitably lost, resulting in inaccurate captions. Therefore, the knowledge generated based on these captions may also be inaccurate. These inaccurate decision contexts may lead to incorrect answers. In extreme cases, if all answer candidates are wrong, our method will fail. Hence, improving the quality of decision contexts is an important direction for future research. Moreover, our method requires multiple calls to the LLM for inference, which allows inference-time scaling up of model capabilities but may take too long. Reducing the running time could be a direction for future research.
Although our work has achieved good results, it also inherits any existing biases of LLMs and their training datasets. Future work can focus on addressing these issues.
Appendix A Experimental Details
We use the instruction-tuned LLM in generation of the knowledge text and the automatic rationale, and the LLM before instruction tuning in generation of the answer and the mechanistic rationale. For the standard version of the LLM, we use a greedy decoding strategy. For the instruction-tuned version of the LLM, we use top-k sampling, where k is set to 50. We set the temperature, the length penalty and the repetition penalty all to 1.0, and set the diversity penalty to 0.
Appendix B Prompt Templates
@@
Table 7: Prompt template for ensemble. We first provide rationales for each answer, then ask the model to select a final answer from the answer candidates based on these rationales.
Appendix C Interactions Between Decision Contexts
To explore the effects of interactions between decision contexts, we conduct additional experiments using Mistral-7B by concatenating different decision contexts in one prompt. The result is 41.7% (44.7% resp.) on A-OKVQA val split (OKVQA test split resp.), which is much lower than 47.5% (49.2% resp.) of our method. Consistent with our observation, Cao and Jiang (2024) also concatenate captions and long-form knowledge in KGenVQA, and the result is 39.1% (45.4% resp.). We hypothesize that concatenating different decision contexts into a single prompt can introduce contradictory or redundant information, making it difficult for the model to effective utilize the information in the context to answer questions.
Appendix D Ensemble of Different Caption-only Strategies
We conduct experiments using three different image captioning models, BLIP (Li et al., 2022), InstructBLIP (Dai et al., 2023) and OFA (Wang et al., 2022) to generate different captions. We replace the short-form and long-form knowledge in our method with the captions generated by InstructBLIP and OFA. The result with Mistral-7B on A-OKVQA val split (OKVQA test split resp.) is 44.5% (46.7% resp.), much lower than 47.5% (49.2% resp.) from our method.
Appendix E Examples
In this section, we display more examples from OK-VQA where DietCoke can help model select the correct answer, as shown in Fig 7.
Understanding why the LLM selects certain answers is challenging for two main reasons. First, as recent research (West et al., 2023) indicates, LLMs exhibit behaviors drastically different from humans and may not fully comprehend the outputs they generate. That is, even if the LLM generates the right answer, it may not be able to explain why it selects a particular answer, or do other things that we expect a human who make the right choice would be able to do. Second, without ground-truth annotations for errors in the rationales, it is difficult to quantitatively assess if LLMs correctly detect illogical rationales or factual errors.
Despite these challenges, in Fig. 5, we analyze the relation between the rationale surface form containing exactly the answer candidate and the candidate being selected by the answer-selecting ensemble LLM. The positive correlation between the two events suggests that the surface form of the rationales plays a role in the decision of the answer-selecting LLM.
In addition, we provide three examples by asking the LLM to explain its choice to illustrate how the LLM selects answers. We use “Can you explain briefly how you select the final answer based on the three answer candidates and corresponding rationales?” as the instruction, and providing the final answer and three answer candidates with their rationales in the prompt. For question (a) in Fig. 7, the output is “1847 is the final answer because it marks the historical introduction of donuts to North America, supported by corresponding rationale.” We speculate that the rationale behind the answer "1847" is more relevant to the question regarding the "first introduced" date, so the model selects "1847" as the final answer.
For question (b), the output is “The final answer ‘Yacht’ was selected because multiple rationales consistently describe a large luxury vessel in the scene, and the term ‘yacht’ aligns more specifically with that description than the more general term ‘boat’.” For question (c), the output is “I selected the final answer ‘8’ because it provides the most precise and direct explanation by identifying the number of sides (8) of an octagonal stop sign.”
We also provide some examples containing wrong rationales in Fig. 8.
Figure 7: Examples from OKVQA where our method can help model select the correct answer. From top to bottom are the caption-only strategy, the short-form knowledge strategy, and the long-form knowledge strategy. Incorrect answers are marked with red and correct answers are in green.
Figure 8: Examples from OKVQA where certain rationales are wrong. The wrong rationales and answers are marked in red, and correct answers are in green.